# Majorization via positivity of Jack and Macdonald polynomial differences

### v1

@@ Abstract.
Majorization inequalities have a long history, going back to Maclaurin and Newton. They were recently studied for several families of symmetric functions, including by Cuttler–Greene–Skandera (2011), Sra (2016), Khare–Tao (2021), McSwiggen–Novak (2022), and Chen–Sahi (2024+) among others. Here we extend the inequalities by these authors to Jack and Macdonald polynomials, and obtain conjectural characterizations of majorization and of weak majorization of the underlying partitions. We prove these characterizations for several cases of partitions, including all partitions with two parts. In fact, we upgrade – and prove in the above cases – the characterization of majorization, to containment of Jack and Macdonald differences lying in the Muirhead semiring.
@@ 1.1. Symmetric functions
In the late 1960s, Jack discovered Jack polynomials, unifying Schur polynomials and zonal polynomials – the latter being related to the representation theory of $GL_{n}(\mathbb{R})$ and multivariate statistics.
@@ 1.2. Symmetric function inequalities
The notion of positivity has an even longer history: it begin with mathematics itself – counting and measuring objects. Let us restrict our attention to inequalities about symmetric functions.
Most families of symmetric polynomials are indexed by integer partitions. An integer partition (of length at most $n$) is an $n$-tuple $\lambda=(\lambda_{1},\dots,\lambda_{n})$ of integers such that $\lambda_{1}\geqslant\cdots\geqslant\lambda_{n}\geqslant 0$. Denote the set of these partitions by $\mathcal{P}_{n}$.
@@ 1.2.1. Containment-type and expansion positivity
We say a symmetric polynomial is Schur positive if it can be written as a sum of Schur polynomials with non-negative (integer) coefficients. For example, by the celebrated Littlewood–Richardson rule, the product of two Schur polynomials is Schur positive: $s_{\lambda}\cdot s_{\mu}=\sum_{\nu}c_{\lambda\mu}^{\nu}s_{\nu}$, where the Littlewood–Richardson coefficients $c_{\lambda\mu}^{\nu}\geqslant 0$. Similarly, the Schur polynomial $s_{\lambda}$ is monomial positive: $s_{\lambda}=\sum_{\mu\mathrel{\succcurlyeq}\lambda}K_{\lambda\mu}m_{\mu}$, where the Kostka numbers $K_{\lambda\mu}\geqslant 0$.
Expansion positivity naturally suggests underlying combinatorial objects or structure. For example, Littlewood–Richardson coefficients have various combinatorial interpretations: Littlewood–Richardson tableaux [27, 18], Berenstein–Zelevinsky patterns [3] and Knutson–Tao honeycombs [13]; and Kostka numbers count the number of semi-standard Young tableaux (SSYT) [27, 18].
For Jack polynomials $P_{\lambda}^{(\alpha)}(x)$, it was conjectured in Macdonald’s book [18, VI.(10.26?)] and proved by Knop and Sahi [12] that their expansions in terms of the monomial basis $(m_{\lambda})$ are positive (and integral, under certain normalizations) in the parameter $\alpha$.
@@
$\binom{\lambda}{\nu}_{\alpha}-\binom{\mu}{\nu}_{\alpha}\in\mathbb{F}_{\geqslant 0},\quad\lambda\supseteq\mu$ (1.2)
@@
The study of generalized binomial coefficients originated in two distinct contexts. In statistics, their zonal case ($\alpha=2$) was first examined by Constantine in the 1960s, while Lascoux [15] explored their Schur case ($\alpha=1$) to compute Chern classes for exterior and symmetric squares of vector bundles (see also [18, p. 47, Example 10]). Seminal contributions emerged in the 1990s through the foundational work of Lassalle [16, 17], Kaneko [9], and Okounkov–Olshanski [22, 21].
@@
Theorem 1.1 ([4, Theorem 6.1]).
The following statements are equivalent:
@@
  2. (2)
    For any fixed $\tau_{0}\in[0,\infty]$, the difference $\frac{P_{\lambda}(x+\bm{1};\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(x+\bm{1};\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}$ is $\tau_{0}$-Jack positive over $\mathbb{R}_{\geqslant 0}$, namely, can be written as an $\mathbb{R}_{\geqslant 0}$-combination of $P_{\nu}(x;\tau_{0})$.
  3. (3)
It is well-known, see [18, Chapters I, VI, VII], that Jack polynomials $P_{\lambda}(x;\tau)$ specialize to many other symmetric polynomials: monomial symmetric polynomials $m_{\lambda}$ when $\tau=0$, zonal polynomials $Z_{\lambda}$ when $\tau=1/2$ (spherical functions for $GL_{n}(\mathbb{R})/O(n)$) and $\tau=2$ (for $GL_{n}(\mathbb{H})/Sp(n)$), Schur polynomials $s_{\lambda}$ when $\tau=1$, and elementary symmetric polynomials $e_{\lambda^{\prime}}$ when $\tau=\infty$ (where $\lambda^{\prime}$ is the conjugate of $\lambda$). Hence, we also have the following characterization, which somewhat surprisingly also holds for power-sums:
Theorem 1.2 ([4, Theorems 6.2, 6.4]).
The following statements are equivalent:
@@ 1.2.2. Majorization-type and evaluation positivity
$\frac{m_{\lambda}(x)}{m_{\lambda}(\bm{1})}\geqslant\frac{m_{\mu}(x)}{m_{\mu}(\bm{1})},\quad x\in(0,\infty)^{n}$ (1.6)
@@
Inspired by the inequalities above (and some others), Cuttler, Greene, and Skandera [5] generalized the inequalities of Newton and Gantmacher to elementary symmetric polynomials and power sums indexed by partitions, and conjectured an inequality for Schur polynomials, which was later proved by Sra [26] (again, for $\lambda,\mu$ real $n$-tuples). We recall all these results.
@@
    ([5, Theorem 3.2], generalized Newton’s inequality) The following difference is positive:
@@
    ([5, Theorem 4.2], generalized Gantmacher’s inequality) The following difference is positive:
@@
    ([5, Conjecture 7.4, Theorem 7.5] and [26], Cuttler–Greene–Skandera and Sra’s inequality) The following difference is positive:
@@
As we will see below – as also in [5, 10, 19, 26] – the more challenging implication is to show that if $\lambda\mathrel{\succcurlyeq}\mu$ then the normalized differences of various symmetric polynomials are non-negative on $0,\infty)^{n}$. An alternate, stronger approach would be to show that these differences “decompose” non-negatively into non-negative quantities, like monomials and differences of monomials as in part (2). This is known for $e_{\lambda^{\prime}}$ and for $p_{\lambda}$, as we now state.
Let $M_{\lambda}\coloneqq m_{\lambda}/m_{\lambda}(\bm{1})$ be the normalized monomial. Recall [5, Section 6] that the Muirhead cone $\mathcal{M}_{C}(\mathcal{C})$ and the Muirhead semiring $\mathcal{M}_{S}(\mathcal{C})$ over a cone $\mathcal{C}$ are defined as follows: $\mathcal{M}_{C}(\mathcal{C})$ consists of $\mathcal{C}$-linear combinations of Muirhead differences $\{\,M_{\lambda}-M_{\mu}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\mathrel{\succcurlyeq}\mu\,\}$, and $\mathcal{M}_{S}(\mathcal{C})$ consists of $\mathcal{C}$-linear combinations of products of functions in $\{\,M_{\lambda}-M_{\mu}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\mathrel{\succcurlyeq}\mu\,\}\cup\{\,M_{\lambda}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\,\}$. In particular, when evaluated at $x\in[0,\infty)^{n}$, functions in $\mathcal{M}_{C}(\mathcal{C})$ and $\mathcal{M}_{S}(\mathcal{C})$ take values in the cone $\mathcal{C}$.
Theorem 1.4 ([5, Theorem 6.1]).
@@
It should be noted that the Muirhead semiring $\mathcal{M}_{S}$ is strictly larger than the Muirhead cone $\mathcal{M}_{C}$. Some power sum differences belong to $\mathcal{M}_{S}\setminus\mathcal{M}_{C}$. For example, when $n=3$, $\lambda=(33)$ and $\mu=(321)$, we have
$\frac{p_{\lambda}}{p_{\lambda}(\bm{1})}-\frac{p_{\mu}}{p_{\mu}(\bm{1})}=\frac{2}{3}M_{(3)}\left(M_{(3)}-M_{(21)}\right)=\frac{2}{9}\left(M_{(6)}-M_{(51)}-M_{(42)}+2M_{(33)}-M_{(321)}\right),$
which is in $\mathcal{M}_{S}$ but not in $\mathcal{M}_{C}$. However, it seems that the differences of elementary symmetric polynomials lie in the Muirhead cone $\mathcal{M}_{C}$.
@@ 1.2.3. Weak majorization-type and evaluation positivity
These inequalities (specifically, the final one in the next result) were first recorded in [10, Theorem 1.14] for Schur polynomials, followed by the other variants that were shown in [4].
Theorem 1.5 ([4, Theorem 6.5]).
The following statements are equivalent:
@@
As noted in [4, Theorem 6.5], these inequalities follow easily from their containment- and majorization- analogues, via the following easy lemma that connects the three partial orders.
Lemma 1.6 (see [4]).
If $\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}\mu$, then there exists some $\nu$ such that $\lambda\supseteq\nu\mathrel{\succcurlyeq}\mu$.
@@ 1.3. Our work
This work was motivated by the goal of extending the above characterizations to Jack and even Macdonald polynomials. We refer the reader to:
  • Theorems 3.1, 3.2, 3.4 and 4.8, connecting (weak) majorization and normalized Jack differences;
@@
  • Theorems 5.6 and 5.8 for (weak) majorization via Macdonald polynomials.
@@ 2.1. Partitions
An (integer) partition is an infinite tuple $\lambda=(\lambda_{1},\lambda_{2},\dots)$ of weakly-decreasing non-negative integers with finitely many nonzero entries. The non-zero $\lambda_{i}$’s are called the parts of $\lambda$. The zeros are usually omitted, and we use the exponential to indicate repeated entries. The length of $\lambda$, denoted by $\ell(\lambda)$, is the number of parts; and the size is the sum of the parts, $|\lambda|=\lambda_{1}+\cdots+\lambda_{\ell(\lambda)}$. For example, $\lambda=(2^{2}1^{3})=(2,2,1,1,1,0,\dots)$ has length 5 and size 7.
@@ 2.2. Symmetric polynomials
Each of $\{\,m_{\lambda}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\in\mathcal{P}_{n}\,\}$, $\{\,e_{\lambda^{\prime}}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\in\mathcal{P}_{n}\,\}$, $\{\,p_{\lambda^{\prime}}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\in\mathcal{P}_{n}\,\}$ and $\{\,s_{\lambda}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\in\mathcal{P}_{n}\,\}$ is a $\mathbb{Q}$-basis of $\Lambda_{n}$ [18, Chapter I].

### v2

@@ Abstract.
Majorization inequalities have a long history, going back to Maclaurin and Newton. They were recently studied for several families of symmetric functions, including by Cuttler–Greene–Skandera (2011), Sra (2016), Khare–Tao (2021), McSwiggen–Novak (2022), and Chen–Sahi (2024+) among others. Here we extend the inequalities by these authors to Jack and Macdonald polynomials, and obtain conjectural characterizations of majorization and of weak majorization of the underlying partitions. We prove these characterizations for two variables. In fact, we upgrade—and prove in the above cases—the characterization of majorization, to containment of Jack and Macdonald differences lying in the Muirhead semiring.
@@ 1.1. Symmetric functions
In the late 1960s, Jack discovered Jack polynomials, unifying Schur polynomials and zonal polynomials—the latter being related to the representation theory of $GL_{n}(\mathbb{R})$ and multivariate statistics.
@@ 1.2. Symmetric function inequalities
The notion of positivity has an even longer history: it begin with mathematics itself—counting and measuring objects. Let us restrict our attention to inequalities about symmetric functions.
Most families of symmetric polynomials are indexed by integer partitions. A partition (of length at most $n$) is an $n$-tuple $\lambda=(\lambda_{1},\dots,\lambda_{n})$ of integers such that $\lambda_{1}\geqslant\cdots\geqslant\lambda_{n}\geqslant 0$. Denote the set of such partitions by $\mathcal{P}_{n}$. Let $|\lambda|\coloneqq\lambda_{1}+\cdots+\lambda_{n}$ be the size of $\lambda$.
@@ 1.2.1. Containment-type and expansion positivity
We say a symmetric polynomial is Schur positive if it can be written as a sum of Schur polynomials with non-negative coefficients. For example, by the celebrated Littlewood–Richardson rule, the product of two Schur polynomials is Schur positive: $s_{\lambda}\cdot s_{\mu}=\sum_{\nu}c_{\lambda\mu}^{\nu}s_{\nu}$, where the Littlewood–Richardson coefficients $c_{\lambda\mu}^{\nu}\geqslant 0$. Similarly, the Schur polynomial $s_{\lambda}$ is monomial positive: $s_{\lambda}=\sum_{\mu\mathrel{\succcurlyeq}\lambda}K_{\lambda\mu}m_{\mu}$, where the Kostka numbers $K_{\lambda\mu}\geqslant 0$.
Expansion positivity naturally suggests underlying combinatorial objects or structure. For example, Littlewood–Richardson coefficients have various combinatorial interpretations: Littlewood–Richardson tableaux [26, 17], Berenstein–Zelevinsky patterns [2] and Knutson–Tao honeycombs [12]; and Kostka numbers count the number of semi-standard Young tableaux [26, 17].
For Jack polynomials $P_{\lambda}^{(\alpha)}(x)$, it was conjectured in Macdonald’s book [17, VI.(10.26?)] and proved by Knop and Sahi [11] that their expansions in terms of the monomial basis $(m_{\lambda})$ are positive (and integral, under certain normalizations) in the parameter $\alpha$.
@@
$\binom{\lambda}{\nu}_{\alpha}-\binom{\mu}{\nu}_{\alpha}\in\mathbb{F}_{\geqslant 0},\quad\lambda\supseteq\mu,$ (1.2)
@@
The study of generalized binomial coefficients originated in two distinct contexts. In statistics, their zonal case ($\alpha=2$) was first examined by Constantine in the 1960s, while Lascoux [14] explored their Schur case ($\alpha=1$) to compute Chern classes for exterior and symmetric squares of vector bundles (see also [17, p. 47, Example 10]). Seminal contributions emerged in the 1990s through the foundational work of Lassalle [15, 16], Kaneko [8], and Okounkov–Olshanski [21, 20].
@@
Theorem 1.1 ([3, Theorem 6.1]).
The following are equivalent for partitions $\lambda,\mu$:
  1. (1)
  2. (2)
    For every $\tau_{0}\in[0,\infty]$, the difference $\frac{P_{\lambda}(x+\bm{1};\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(x+\bm{1};\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}$ is $\tau_{0}$-Jack positive over $\mathbb{R}_{\geqslant 0}$, namely, can be written as an $\mathbb{R}_{\geqslant 0}$-combination of $P_{\nu}(x;\tau_{0})$.
  3. (3)
    For some $\tau_{0}\in[0,\infty]$, the difference $\frac{P_{\lambda}(x+\bm{1};\tau_{0})}{P_{\lambda}(\bm{1};\tau_{0})}-\frac{P_{\mu}(x+\bm{1};\tau_{0})}{P_{\mu}(\bm{1};\tau_{0})}$ is $\tau_{0}$-Jack positive over $\mathbb{R}_{\geqslant 0}$, namely, can be written as an $\mathbb{R}_{\geqslant 0}$-combination of $P_{\nu}(x;\tau_{0})$.
@@
It is well-known, see [17, Chapters I, VI, VII], that Jack polynomials $P_{\lambda}(x;\tau)$ specialize to many other symmetric polynomials: monomial symmetric polynomials $m_{\lambda}$ when $\tau=0$, zonal polynomials $Z_{\lambda}$ when $\tau=1/2$ (spherical functions for $GL_{n}(\mathbb{R})/O_{n}$) and $\tau=2$ (for $GL_{n}(\mathbb{H})/Sp_{n}$), Schur polynomials $s_{\lambda}$ when $\tau=1$, and elementary symmetric polynomials $e_{\lambda^{\prime}}$ when $\tau=\infty$ (where $\lambda^{\prime}$ is the conjugate of $\lambda$). Hence, we also have the following characterization, which somewhat surprisingly also holds for power sums:
Theorem 1.2 ([3, Theorems 6.2, 6.3]).
The following are equivalent for partitions $\lambda,\mu$:
@@ 1.2.2. Majorization-type and evaluation positivity
$\frac{m_{\lambda}(x)}{m_{\lambda}(\bm{1})}\geqslant\frac{m_{\mu}(x)}{m_{\mu}(\bm{1})},\quad x\in[0,\infty)^{n}$ (1.6)
@@
Cuttler, Greene, and Skandera [4] generalized the inequalities of Newton and Gantmacher to elementary symmetric polynomials and power sums indexed by partitions, and conjectured an inequality for Schur polynomials, which was later proved by Sra [25] (again, for $\lambda,\mu$ real $n$-tuples). We recall all these results.
@@
    ([4, Theorem 3.2], generalized Newton’s inequality) The following difference is positive:
@@
    ([4, Theorem 4.2], generalized Gantmacher’s inequality) The following difference is positive:
@@
    ([4, Conjecture 7.4, Theorem 7.5] and [25], Cuttler–Greene–Skandera and Sra’s inequality) The following difference is positive:
@@
As we will see below—as also in [4, 9, 18, 25]—the more challenging implication is to show that if $\lambda\mathrel{\succcurlyeq}\mu$ then the normalized differences of various symmetric polynomials are non-negative on $0,\infty)^{n}$. An alternate, stronger approach would be to show that these differences “decompose” non-negatively into non-negative quantities, like monomials and differences of monomials as in Muirhead’s inequality. This is known for $e_{\lambda^{\prime}}$ and for $p_{\lambda}$, as we now state.
Let $M_{\lambda}\coloneqq m_{\lambda}/m_{\lambda}(\bm{1})$ be the normalized monomial. Recall [4, Section 6] that the Muirhead cone $\mathcal{M}_{C}(\mathcal{C})$ and the Muirhead semiring $\mathcal{M}_{S}(\mathcal{C})$ over a cone $\mathcal{C}$ are defined as follows: $\mathcal{M}_{C}(\mathcal{C})$ consists of $\mathcal{C}$-linear combinations of Muirhead differences $\{\,M_{\lambda}-M_{\mu}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\mathrel{\succcurlyeq}\mu\,\}$, and $\mathcal{M}_{S}(\mathcal{C})$ consists of $\mathcal{C}$-linear combinations of products of functions in $\{\,M_{\lambda}-M_{\mu}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\mathrel{\succcurlyeq}\mu\,\}\cup\{\,M_{\lambda}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\,\}$. In particular, when evaluated at $x\in[0,\infty)^{n}$, functions in $\mathcal{M}_{C}(\mathcal{C})$ and $\mathcal{M}_{S}(\mathcal{C})$ take values in the cone $\mathcal{C}$.
Theorem 1.4 ([4, Theorem 6.1, Corollary 6.2]).
@@
It should be noted that the Muirhead semiring $\mathcal{M}_{S}$ is strictly larger than the Muirhead cone $\mathcal{M}_{C}$; see Section 4.1 below.
@@ 1.2.3. Weak majorization-type and evaluation positivity
These inequalities (specifically, the final one in the next result) were first recorded in [9, Theorem 1.14] for Schur polynomials, followed by the other variants that were shown in [3].
Theorem 1.5 ([3, Theorem 6.7]).
The following are equivalent for partitions $\lambda,\mu$:
@@
As noted in [3], these inequalities follow easily from their containment- and majorization- analogues, via the following easy lemma that connects the three partial orders.
Lemma 1.6 ([3, Lemma 6.6]).
$\lambda\mathrel{\succcurlyeq_{\mathrm{w}}}\mu$ if and only if there exists some $\nu$ such that $\lambda\supseteq\nu\mathrel{\succcurlyeq}\mu$.
@@ 1.3. Our work
As Jack polynomials unify Schur, monomial, and elementary symmetric polynomials, it is a pleasing fact that Theorem 1.1 holds, unifying several related but disparate positive-coefficient expansion-characterizations of containment in Theorem 1.2. In the same vein, it is natural to ask if Theorems 1.3 and 1.5 can also be extended to Jack polynomials for all parameters, or more strongly, to Jack-positivity in the sense of Theorem 1.1(3).
The goal of this work is to achieve the above unifications in several cases. We also go beyond Jack polynomials, to Macdonald polynomials with parameters $q,t\in(0,1)$. We refer the reader to:
  • Theorems 3.4, 3.2 and 4.9, connecting (weak) majorization and normalized Jack differences;
@@
  • Theorem 5.5 for majorization via Macdonald polynomials.
Our results also fit into—and contribute to—the framework of duality discussed in [3]: given a partial order $\mathrel{\succcurlyeq}_{1}$ on symmetric polynomials and a basis $\{b_{\lambda}(x)\}$ of them, as well as a partial order $\mathrel{\succcurlyeq}_{2}$ on partitions, we say that $\mathrel{\succcurlyeq}_{1}$ and $\mathrel{\succcurlyeq}_{2}$ are dual for the family $b_{\lambda}$ if the following holds:
$b_{\lambda}\mathrel{\succcurlyeq}_{1}b_{\mu}\ \text{if and only if}\ \lambda\mathrel{\succcurlyeq}_{2}\mu.$ (1.16)
In [3], two of us showed that the containment order $\mathrel{\succcurlyeq}_{2}$ is dual to “Jack expansion positivity” (wherein $p(x)\mathrel{\succcurlyeq}_{1}q(x)$ means pqp-q is a positive combination of Jack polynomials—for a specific choice of parameter or for an indeterminate parameter). In this spirit, Theorem 1.3 asserts that majorization, or the dominance order $\mathrel{\succcurlyeq}$ on partitions (which we take to be $\mathrel{\succcurlyeq}_{2}$) is dual to “evaluation positivity” on the positive orthant: $p\mathrel{\succcurlyeq}_{1}q$ if $p(x)\geqslant q(x)$ for all $x\in(0,\infty)^{n}$—this is for normalized Schur, or monomial, or other specific families of symmetric functions.
Similarly, Theorem 1.5 shows the duality between the weak dominance order $\mathrel{\succcurlyeq}_{2}$ (which is now $\mathrel{\succcurlyeq_{\mathrm{w}}}$) and evaluation positivity for the shifted normalized bases $f_{\lambda}(x+{\bf 1})$ (with the same $f$ as previously) on the positive orthant. It is our goal in this work to extend both of these dualities to Jack polynomials with arbitrary parameter—and even more strongly, to Macdonald polynomials.
@@ 2.1. Partitions
A partition is an infinite tuple $\lambda=(\lambda_{1},\lambda_{2},\dots)$ of weakly-decreasing non-negative integers with finitely many nonzero entries. The non-zero $\lambda_{i}$’s are called the parts of $\lambda$. The zeros are usually omitted, and we use the exponential to indicate repeated entries. The length of $\lambda$, denoted by $\ell(\lambda)$, is the number of parts; and the size is the sum of the parts, $|\lambda|=\lambda_{1}+\cdots+\lambda_{\ell(\lambda)}$. For example, $\lambda=(2^{2}1^{3})=(2,2,1,1,1,0,\dots)$ has length 5 and size 7.
@@ 2.2. Symmetric polynomials
Each of $\{\,m_{\lambda}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\in\mathcal{P}_{n}\,\}$, $\{\,e_{\lambda^{\prime}}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\in\mathcal{P}_{n}\,\}$, $\{\,p_{\lambda^{\prime}}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\in\mathcal{P}_{n}\,\}$ and $\{\,s_{\lambda}\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}\lambda\in\mathcal{P}_{n}\,\}$ is a $\mathbb{Q}$-basis of $\Lambda_{n}$ [17, Chapter I].